OpenStudy (anonymous):

6.8 * 〖10〗^7 divided by 5.3 *〖10〗^(-3) equals

OpenStudy (anonymous):

12830.18867925 i have no clue if i am wrong or not, but thats what i get

OpenStudy (anonymous):

are you doing algebra?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

\[6.8\times 10^7\times 5.3\times 10^3\] is a start

OpenStudy (anonymous):

then compute \[6.8\times 5.3\times 10^{7+3}\]